OpenStudy (anonymous):

In a direct variation, y is 4 when x is twelve. What is y when x equals six? negative two 2 3 eighteen

OpenStudy (shamil98):

Formula for direct variation is: \[y = kx\] You are given this: y = 4 when x = 12 so we need to find what the variable k is (k is a constant). \[k = \frac{ y }{ x } = \frac{ 4 }{ 12 } = \frac{ 1 }{ 3 }\] so you need to find what y is when x = 6, just plug in your values of k and x into \[y = kx \] to find the value of y.

OpenStudy (anonymous):

how do i know what is

OpenStudy (anonymous):

what k is

OpenStudy (shamil98):

I just explained it.

OpenStudy (anonymous):

is it 18

OpenStudy (anonymous):

I.... Idekrn

OpenStudy (anonymous):

? ¿